OpenStudy (anonymous):

The region in the first quadrant bounded by the x-axis, the line x = π, and the curve y = sin(sin(x)) is rotated about the x-axis. What is the volume of the generated solid?

OpenStudy (anonymous):

https://gyazo.com/9ccb696779c5a1339d9c084cfe54d9e5 @freckles

OpenStudy (anonymous):

@Zarkon

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi} \pi[\sin(\sin (x))]^2 dx \]

OpenStudy (anonymous):

\[A(x)=\pi r^2\]

OpenStudy (anonymous):

I am getting 5.435 x 10^-5 Which is obviously wrong

OpenStudy (anonymous):

@zepdrix Are you aware of what must be done for this question?

zepdrix (zepdrix):

Sec, lemme see if I can work it out :)

OpenStudy (anonymous):

Ok thanks

zepdrix (zepdrix):

Hmm your integral looks correct :o Where did you get the numerical value from? I don't think this integral has a nice closed-form solution. So I guess you have to use some type of algebra calculator or something? Wolfram seems to give us the correct answer: https://www.wolframalpha.com/input/?i=integral+of+pi+sin%5E2%28sinx%29+from+x%3D0+to+x%3Dpi

OpenStudy (anonymous):

I used my calculator to save me some time.

zepdrix (zepdrix):

Ya you would have to do that for this problem :) Maybe some trouble inputting it with your brackets and all that? Cause your setup looks good.

OpenStudy (anonymous):

I entered it as above

OpenStudy (anonymous):

Maybe the way you entered it into wolfram is more appropriate, let me try.

zepdrix (zepdrix):

Wolfram is pretty good at interpreting words

OpenStudy (anonymous):

Does the mode matter?

zepdrix (zepdrix):

The boundaries of your integral are radian measure. Hmm, so yes, maybe.

OpenStudy (anonymous):

I'm getting a data type error. Is there a way to insert sin^2?

zepdrix (zepdrix):

sin^2(sin(x)) is the same as (sin(sin(x)))^2 so maybe try that second way oh, I guess that's what you had in the first place though.

OpenStudy (anonymous):

Yeah, I had that. Tells me wrong value or variable type.

zepdrix (zepdrix):

Notice that Wolfram's exact solution involves the Bessel function of the first kind. So maybe your calculator can't handle fancy stuff like this one. I'm not sure :\

OpenStudy (anonymous):

LOL I can't use wolfram when taking my exam doe, they proctor my online course exam at school, when my own teacher tells me I can cheat.

zepdrix (zepdrix):

hmm true >.<

OpenStudy (anonymous):

I'm just gonna use it and tell em my calculator is acting up so I'm using an online one.

OpenStudy (anonymous):

Thanks man

zepdrix (zepdrix):

ya np sorry we couldn't figure out the calculator :\ maybe ask teacher for assistance

OpenStudy (anonymous):

Quick question if you don't mind.

OpenStudy (anonymous):

and don't worry about it

OpenStudy (anonymous):

I have a questions extremely similar to this one but has an equilateral triangle instead. Would I follow same steps but use the area formula of the triangle?

OpenStudy (anonymous):

Oh wait, that one is cross sections.....it's different

zepdrix (zepdrix):

maybe post the question? :) If I can look at it, it might refresh my memory

OpenStudy (anonymous):

Ok, I'll tag you soon.